OpenStudy (anonymous):

15. A restaurant has one type of milk that has 2% fat and another that has 7% fat. how many quarts of each type does the restaurant need to make 40 quarts of a milk mixture that is 5% fat? a. 24 quarts of the 2% milk and 16 quarts of the 7% milk b. 16 quarts of the 2% milk and 24 quarts of the 7% milk c. 16 quarts of the 2% milk and 16 quarts of the 7 milk d. 24 quarts of the 2% milk and 24 quarts of the 7% milk

OpenStudy (anonymous):

oh boy a mixture problem ...

OpenStudy (anonymous):

@dtan5457

OpenStudy (anonymous):

yeah o_o

OpenStudy (anonymous):

lets call the amount of 7% milk x so the amount of 2% milk is \(40-x\) because the total is 40

OpenStudy (anonymous):

\(5\% \) of \(40\) is \(2\) right?

OpenStudy (anonymous):

that means \[0.07x+0.02\times (40-x)=2\] or to make it easier \[7x+2(40-x)=200\]

OpenStudy (anonymous):

O_O I-I think

OpenStudy (dtan5457):

well this is how i did it.. x+y=40 2x+7y/40=5 2x+7y=200 x=40-y plug in to get 80-2y+7y=200 5y=200 y=24 x=16 b. 16 quarts of the 2% milk and 24 quarts of the 7% milk

OpenStudy (anonymous):

yeah it is because ten percent of forty is four, so five percent is two

OpenStudy (dtan5457):

@satellite73 please check my work

OpenStudy (dtan5457):

wait nvm i already screwed up 5y=120

OpenStudy (dtan5457):

not 200..

OpenStudy (anonymous):

o.o

OpenStudy (anonymous):

at some point you are going to have to solve \[7x+2(40-x)=200\]

OpenStudy (anonymous):

wow wow wow wow wow wow wow wow wow wow. now I gotta read all of this

OpenStudy (slinley):

yes! @Mimi_Styles TO understand IT!!!

OpenStudy (dtan5457):

if i just solve for y for 7% milk, isn't that enough?

OpenStudy (anonymous):

ok we can start again if you like and do it slow

OpenStudy (anonymous):

these problems are all more or less the same do one or two and you can do them all

OpenStudy (anonymous):

<____>

OpenStudy (anonymous):

i don't know if that is a yes or a no

OpenStudy (dtan5457):

i just wanna know if i my answer is correct so i can rest in peace

OpenStudy (anonymous):

I made this face and it's called the confused face

OpenStudy (anonymous):

ok first off, do you know that \(7\%\) means \(.07\)

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

and do you know that 7 percent of a number means that number times \(.07\) ?

OpenStudy (anonymous):

if you do not, feel free to say "i know it now"

OpenStudy (anonymous):

I know now... ._.

OpenStudy (anonymous):

ok so for example 7% of 50 is \(.07\times 50\) and 7% of \(x\) is \(.07x\)

OpenStudy (anonymous):

o.o

OpenStudy (anonymous):

good to here?

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

if we are gong to solve one of these, we have to have a variable some people actually like two variables, i think that is more confusing, one will do and we have to say what it is i say "let \(x\) be the amount of 7% milk"

OpenStudy (anonymous):

then in terms of \(x\) we know that the total number of gallons is 40 if \(x\) of them are 7% milk the rest, the 2% milk is \(40-x\) for example if there were 15 gallons of 7% milk there would have to be \(40-15=25\) gallon of 2% milk

OpenStudy (anonymous):

since we now know that 7% of x is \(.07x\) and also that 2% of \(40-x\) is \(.02(40-x)\) we can make an equation

OpenStudy (anonymous):

the total amount of fat or whatever is 5% of 40, which we decided is 2 that means \[.07x+.02(40-x)=2\]

OpenStudy (anonymous):

o.o ohh

OpenStudy (anonymous):

now that is somewhat annoying to solve

OpenStudy (anonymous):

but not too bad if you multiply both sides by 100 and solve \[7x+2(40-x)=200\]

OpenStudy (anonymous):

i am sure you can do that one once you get it, you will have the amount of 7% milk

OpenStudy (anonymous):

uhhh D:

OpenStudy (anonymous):

wait wait *lost*

OpenStudy (anonymous):

what step

OpenStudy (anonymous):

"but not too bad if you multiply both sides by 100 and solve 7x+2(40−x)=200 " <----

OpenStudy (anonymous):

that one.

OpenStudy (anonymous):

\[.07x+.02(40-x)=2\] is what needs to be solved and you can do it with the decimals in place if you like

OpenStudy (anonymous):

but multiplying by 100 moved the decimal over two places, i.e. gets rid of them

OpenStudy (anonymous):

if you multiply by 100 you have to multiply both decimals by 100 to get rid of the decimal, but you also have to multiply 2 by 100 to get 200 that is why \[.07x+.02(40-x)=2\] is the same as \[7x+2(40-x)=200\]

OpenStudy (anonymous):

ohh~

OpenStudy (anonymous):

ohhh! ^_^

OpenStudy (anonymous):

so what im doing is...multiplying each sides by 100 ?

OpenStudy (anonymous):

yes to get from the equation with decimals to the equation with out them you still have to solve it though

OpenStudy (anonymous):

0.02 x + 0.07(40 -x ) =40(0.05) 2x + 280-7x = 200 -5x =-80 x = 16 fat = 40-16 = 24

OpenStudy (anonymous):

ok looks reasonable

OpenStudy (anonymous):

i am not sure how you got that one, but it is correct also

OpenStudy (anonymous):

o.o

OpenStudy (dtan5457):

wait so 16 2% and 24 7%?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@Mimi_Styles switched them

OpenStudy (anonymous):

\[0.02 x + 0.07(40 -x ) =40(0.05) \] has \(x\) as the 2% milk

OpenStudy (anonymous):

btw "whole milk" is about \(3.25\%\) or so not 5% and certainly not 7%

OpenStudy (anonymous):

oooohhh